Hello, i need help solving this problem. i am not able to put a graph on here but i will write the needed info.
Determine the equation of the line
y =
(simplify you answer, type answer in slope-intercept form)
there is a graph that has the points (0,5) (2,4)(-2,6)

Answer by Alan3354(69443)

Determine the equation of the line
y =
(simplify you answer, type answer in slope-intercept form)
there is a graph that has the points (0,5) (2,4)(-2,6)
--------------
Use any 2 of the points to find m, the slope
m = diffy/diffx = (5-4)/(0-2) = -1/2
(0,5) is the y-intercept since x = 0
----
--> eqn is y = mx + b
y = (-1/2)x + 5